# GW231123: a product of successive mergers from $\sim 10$ stellar-mass black holes

### v1

@@ 3.2 How many stars are needed to generate GW231123?
Figure 3: Comparison of GW recoil kick velocities (left panel) and final remnant spins (right panel) for mergers with BHs of different generations. The left panel shows the kick velocity distributions, indicating that hierarchical mergers involving higher-generation component BHs produce significantly larger kicks than 1G+1G mergers. The right panel shows the corresponding distributions of the remnants’ final spin magnitudes, with shaded areas comparing the results of GW231123. The 1G+1G, nG+1G, and nG+nG BBHs samples are drawn from the posterior population distribution in Li et al. (2024b).
4 Conclusions and Discussion
GW231123 is the most massive BBH merger detected by LVK to date (Abbott et al., 2023a), making it one of the most intriguing events for tests of astrophysical and fundamental physics (The LIGO Scientific Collaboration et al., 2025).
In this work, we analyzed the data of GW231123 under a variety of population-informed priors corresponding to different formation and evolutionary scenarios. Our results show that neither component of GW231123 is consistent with being non-spinning, effectively ruling out a PBH origin (since primordial BHs are expected to have essentially zero spin in certain models; e.g., Chiba & Yokoyama 2017). A low-spin origin for the GW231123 components is also disfavored–especially for the primary BH. Notably, the scenario in which the secondary BH belongs to the low-spin, low-mass subpopulation (i.e. is a first-generation BH, as defined in Li et al. 2024b) cannot be entirely excluded if the primary BH is highly spinning. This suggests that GW231123 could be a “straddling” binary straddling the PIMG, although it likely did not form from the direct collapse of two massive stars (as was once proposed for GW190521; Fishbach & Holz 2020).
Overall, both components of GW231123 are best described by the high-spin subpopulation, consistent with the higher-generation BHs (Gerosa & Fishbach, 2021; Li et al., 2024b). Moreover, the component masses of GW231123 are exceptionally large compared to those in previous BBH observations–particularly in the case of the primary. In fact, the primary’s mass (and might the secondary’s as well) is more than twice the commonly assumed lower bound of the PIMG. Besides, the spin magnitudes of the two components tend to be more consistent with the remnants of mergers with higher-generation BHs, as shown in Figure 3 (right panel). Therefore, it is more likely that both components of GW231123 have undergone successive mergers with several first-generation (stellar-collapse) BHs.
@@ 4 Conclusions and Discussion
Using the distribution of first-generation BH masses from dynamical formation channels identified in GWTC-3 by Li et al. (2024a) (see also Wang et al. 2022; Ray et al. 2024), we estimate that the primary (secondary) BH in GW231123 could be built up by the merger of $\sim 6$ ($\sim 4$) first-generation BHs. The assembly of such a massive system via hierarchical mergers requires that two key conditions be met. First, earlier merger remnants must be retained in the same environment to undergo subsequent mergers; this means the host’s escape velocity must exceed the gravitational recoil (kick) velocity of each merger remnant. Hierarchical BH mergers impart substantially larger kicks than first-generation mergers (see the left panel of Figure 3), so only environments with sufficiently deep potential wells–such as nuclear star clusters or AGN disks–could keep the GW231123 progenitors bound after each merger. Second, the binary orbits must be harden efficiently so that multiple mergers can occur within a reasonable timescale. In dense stellar systems, dynamical interactions (e.g., binary-single encounters) help harden binaries, but an even more efficient mechanism is likely needed for rapid, repeated mergers. Gas dynamical friction in AGN disks can rapidly shrink binary orbits and shorten merger timescales (Yang et al., 2019), and simulations suggest that this gas-assisted hardening may dominate the upper limit on BBH merger masses (Vaccaro et al., 2024; Xue et al., 2025). Therefore, among known environments, AGN disks are the more plausible sites for producing BBH mergers as massive as GW231123.
@@
If AGN-disk channels indeed dominate the formation of GW231123-like events, then BBH merger masses up to $\sim\mathcal{O}(10^{4})M_{\odot}$ might be achievable (Vaccaro et al., 2024), which are well beyond the detection range of the current generation of ground-based detectors (LIGO Scientific Collaboration et al., 2015; Acernese et al., 2015; Akutsu et al., 2018). The next generation of gravitational-wave observatories, including planned space-based detectors (Luo et al., 2016; Hu & Wu, 2017; Amaro-Seoane et al., 2017; Abbott et al., 2017; Punturo et al., 2010), will be required to observe such extremely massive mergers, offering a more complete view of the BBH mass spectrum and evolutionary pathways. Beyond standard astrophysical channels, a few exotic formation scenarios have been proposed. For example, “dark stars” (hypothetical early-universe stars powered by dark matter annihilation) could form IMBHs; however, the expected masses of such remnants are far larger than those of GW231123 (Lei et al., 2025a). Another proposal involves a cosmological coupling mechanism that gradually increases BH masses, which in principle might produce BHs as massive as GW231123 (Croker et al., 2021). This cosmological coupling hypothesis, however, has been ruled out by very recent JWST observations (Lei et al., 2025b). Ultimately, a comprehensive population analysis of BBH mergers will be necessary to fully elucidate the origins and formation channels of GW231123 and similar exceptional events (e.g., Li et al., 2021, 2022a; Abbott et al., 2023b; Li et al., 2024b; Guo et al., 2024; Alvarez-Lopez et al., 2025).
Acknowledgments
@@
This work is supported by the National Natural Science Foundation of China (No. 12233011, No. 12203101, and No. 12303056), the General Fund (No. 2024M753495) of the China Postdoctoral Science Foundation, and the Priority Research Program of the Chinese Academy of Sciences (No. XDB0550400). This research has made use of data and software obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes.
Appendix A Bayesian Inference
We analyze an 8-s segment of strain data for GW231123, spanning from $t_{\rm c}-6{\rm s}$ to $t_{\rm c}+2{\rm s}$ (where $t_{\rm c}$ is the geocentric GPS time of coalescence), which are available from the Gravitational Wave Open Science Center Catalog [^4] Download from https://doi.org/10.5281/zenodo.15832843. For the noise power spectral density (PSD), we use the pre-estimated PSD files released by The LIGO Scientific Collaboration et al. (2025).
Our primary waveform model is the precessing multipolar approximant IMRPhenomXPHM (Khan et al., 2019; Hannam et al., 2014; Pratten et al., 2021), which includes subdominant harmonics. As a cross‐check, we also perform inference with IMRPhenomXO4a (Thompson et al., 2024), which was shown to exhibit a distinct secondary mode in the component‐mass posterior in The LIGO Scientific Collaboration et al. (2025). The corresponding results for this waveform model are presented in Table 2.
[…]
where we take $f_{\rm min}=20$ Hz and $f_{\rm max}=488$ Hz following The LIGO Scientific Collaboration et al. (2025). Except for the population-informed priors introduce in Section 2, all other priors follow those of The LIGO Scientific Collaboration et al. (2025). We sample the posterior distributions using the Bilby (Ashton et al., 2019b) framework with PyMultiNest (Buchner, 2016) sampler, and verify results with Dynesty (Speagle, 2020).
Table 2: Bayes factors obtained with IMRPhenomXO4a
$\ln\mathcal{B}^{P_{0}}_{x}$ $a_{1}\sim P_{\rm NS}(a)$, $a_{1}\sim P_{\rm LS}(a)$, $a_{1}\sim P_{\rm HS}(a)$, $a_{1}\sim P_{\rm LS}(a)$,
$m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\phi}(m)$ $m_{1}\sim P_{\rm IM}(m)$
$a_{2}\sim P_{\rm NS}(a)$,$m_{2}\sim P_{\phi}(m)$ 40.1 11.6 -0.5 11.1
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\phi}(m)$ 22.9 10.2 0.3 10.3
$a_{2}\sim P_{\rm HS}(a)$,$m_{2}\sim P_{\phi}(m)$ 8.0 8.3 -1.2 7.5
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\rm IM}(m)$ 23.1 9.0 -2.2 12.4
$a_{2}\sim P_{\rm LS}(a)$,$m_{2}\sim P_{\rm SM}(m)$ 19.9 9.3 -3.2 9.9
Note: The Bayes factors with $\ln\mathcal{B}^{P_{0}}_{x}<0$ ($\ln\mathcal{B}^{P_{0}}_{x}<2.3$) are marked in blue (orange), denoting that
the $x$ prior is favored (not ruled out), comparing to the default prior $P_{0}$.
Appendix B Self-consistency check for BH generations
To validate our choice of the first-generation mass distribution in dynamical formation channels, we compare the mass distribution of remnants made by multiple ($n$) first-generation BHs with those of subpopulations revealed in Li et al. (2024a) and Li et al. (2024b). As shown in the top panel of Figure 4, if we assume the first-generation BHs are drawn from the potential dynamical 1G subpopulation (green shaded area) revealed in Li et al. (2024a), remnants from $\gtrsim 2$ progenitors can well reproduce the higher-generation subpopulation (blue shaded area). Additionally, the higher-generation BHs are dominated by the remnants made by 2 first-generation BHs, which is natural for the scenario in hierarchical mergers. However, if we assume the first-generation BHs are drawn from the low-spin subpopulation (orange shaded area) as revealed in Li et al. (2024b), then the mass distribution of remnants made by 2 first-generation BHs would peak at $\sim 20M_{\odot}$, which is inconsistent with the distribution of the higher-generation subpopulation (blue shaded area), see the bottom panel of Figure 4. This discrepancy further supports our adoption of the dynamical first-generation subpopulation for modeling hierarchical growth.

### v2

@@ II.2 Likely higher than second-generation BHs
Figure 2: Comparison of final remnant spins (left panel) and GW recoil kick velocities(right panel) for mergers with BHs of different generations. The left panel shows the corresponding distributions of the remnants’ final spin magnitudes, with shaded areas comparing the component spin magnitudes of GW231123, reweighed with HS-HS priors. The right panel shows the kick velocity distributions, indicating that hierarchical mergers involving higher-generation component BHs produce significantly larger kicks than 1G+1G mergers. The 1G+1G, nG+1G, and nG+nG BBHs samples are drawn from the posterior population distribution in Li et al. (2024b).
Figure 3: Mass distributions of merger remnants formed from first-generation BHs in dynamical formation channels, compared to the component masses of GW231123, reweighed with HS-HS prior. The first-generation BH mass distribution is taken from Li et al. (2024a). The pink and grey shaded regions indicate GW231123’s primary and secondary mass distributions, which align with mass distribution of remnants assembled from $\sim 6$ (green) and $\sim 4$ (blue) first-generation BHs, respectively. A self-consistency check comparing these model predictions to the observed higher-generation BH population is provided in Appendix C.2 (see Figure 5).
As we have found that both components of GW231123 are more likely to be BHs beyond the second generation, it is worth investigating how many stellar-formed BHs constitute them. For this purpose, we have to specify the underlying first-generation BHs that contribute to the successive mergers. Theoretical predictions and population analysis suggest that isolated binary evolution channels contribute to the LVK’s BBH mergers (see e.g. Zevin et al., 2021; Wang et al., 2022; Godfrey et al., 2023; Li et al., 2024a), which are not expected to participate in subsequent hierarchical mergers (Gerosa & Fishbach, 2021). Only the BHs in dynamical formation channels may undergo hierarchical mergers, and potentially contribute to the GW231123-like events. Some population analyses revealed a subpopulation of BHs, which dominates the mass range $\sim[15,40]~M_{\odot}$ (see e.g. Wang et al., 2022; Ray et al., 2024; Li et al., 2024a) [^5] See also Banagiri et al. (2025) for analysis on GWTC-4, though their interpretation is different from us.. This subpopulation is consistent with dynamical formation channels predicted by simulations (see e.g., Antonini & Gieles, 2020; Torniamenti et al., 2022; Antonini et al., 2023; Ford & McKernan, 2025). We note that the results are slightly different in Godfrey et al. (2023), which may be caused by the fact that the models in Godfrey et al. (2023) have only two subpopulations when fitting the data that includes high-spin events, whereas models in Li et al. (2024a) consider three potential subpopulations. In general, there will be at least three subpopulations of BBHs: the first is for the field channels, the second is for the first-generation BBHs from dynamical channels, and the third is for the hierarchical mergers. We note that there are two modes in the posterior distribution of analysis from Godfrey et al. (2023) (see their Figure 6), one mode (taking 76% of samples) may have mixed BBHs from field channels and first-generation BBHs from dynamical channels (Dyn-1G) together, the other mode (taking 24% of samples) may have mixed Dyn-1G BBHs with hierarchical mergers. Therefore, in estimating how many stellar-mass BHs are required to build GW231123’s components, we adopt the possible mass distribution of the Dyn-1G BHs identified by Li et al. (2024a) [^6] Data and codes are available at GitHub: Field V.S. Dyanmical(JackLee0214, 2025). Though the inferred mass function remains largely uncertain, it is broadly consistent with the prediction from some simulations (e.g., Antonini & Gieles, 2020; Torniamenti et al., 2022; Antonini et al., 2023; Ford & McKernan, 2025). Please see Appendix C.1 for the details about generating the mass functions of higher-generation BHs.
To validate our choice of mass function for the Dyn-1G BHs, we have also made a self-consistency check, see Appendix C.2. We compare the mass functions of remnants (made by $n$ Dyn-1G BHs) with the subpopulation of potentially [^7] Such a subpopulation is quite solid now, as was revealed with many different approaches in GWTC-3(e.g. Wang et al., 2022; Pierra et al., 2024b; Antonini et al., 2025c), and recently confirmed with GWTC-4 (e.g. Antonini et al., 2025b; Banagiri et al., 2025; Li et al., 2025b; Wang et al., 2025b). higher-generation BHs observed in GWTC-3 (Li et al., 2024b, a). We find that the distribution of higher-generation BH masses is consistent with remnants produced by the mergers with $\gtrsim 2$ first-generation BHs, supporting the idea that many of the heavy BHs in GWTC-3 (and GW231123 itself) are products of hierarchical growth.
Figure 3 shows the mass distribution of hierarchical merger remnants produced by BHs drawn from this subpopulation, compared against the observed masses of GW231123. From these distributions, we infer that producing a BH of GW231123’s primary mass would require merging roughly 6 first-generation BHs, while the secondary could be built from about 4 first-generation BHs (implying a total equivalent of $\sim 10$ stellar-mass BHs to assemble the system’s $\sim 250~M_{\odot}$ total mass).
If only considering the masses of GW231123, the primary component could also arise from a direct merger between a $\sim 125M_{\odot}$ stellar-collapse IMBH above the PIMG and a $\sim 35M_{\odot}$ BH, which would shorten the merger chain to generate GW231123. However, the predicted final spin in such a scenario is significantly lower than the measured primary spin of GW231123, see right panel of Figure 2. Note that we assume the first-generation BHs are slowly spinning as was found in Li et al. (2024b) (though see also Popa & de Mink, 2025; Croon et al., 2025). Consequently, it is more plausible that both components of GW231123 were assembled through successive mergers of first-generation BHs below the PIMG.
III Conclusions and Discussion
GW231123 is the most massive BBH merger detected by LVK to date (Abbott et al., 2023a), making it one of the most intriguing events for tests of astrophysical and fundamental physics (Abac et al., 2025; Wang et al., 2025a; Tang et al., 2026; Caputo et al., 2025). We establish an association between GW231123 and the populations inferred from previous observations, exploring the formation and evolution of GW231123 in light of BBH subpopulation studies that were carried out with GWTC-3 (Li et al., 2024b) (see also e.g., Wang et al., 2022; Pierra et al., 2024a; Antonini et al., 2025c; Li et al., 2025a; Antonini et al., 2025a, for similar results), with further confirmation from GWTC-4 analyses (e.g., Tong et al., 2025; Antonini et al., 2025b; Banagiri et al., 2025; Li et al., 2025b; Wang et al., 2025b). We show that both components of GW231123 are best described by the high-spin subpopulation, which are natural for the higher-generation BHs (Gerosa & Fishbach, 2021). Besides, the spin magnitudes of the two components tend to be more consistent with the remnants of mergers with higher-generation BHs, as shown in Figure 2 (right panel). Therefore, it is more likely that both components of GW231123 have undergone successive mergers with several first-generation (stellar-collapse) BHs.
@@ III Conclusions and Discussion
Using the distribution of first-generation BH masses from dynamical formation channels identified in GWTC-3 by Li et al. (2024a) (see also Wang et al. 2022; Ray et al. 2024), we estimate that the primary (secondary) BH in GW231123 could be built up by the merger of $\sim 6$ ($\sim 4$) first-generation BHs. The assembly of such a massive system via hierarchical mergers requires that two key conditions be met. First, earlier merger remnants must be retained in the same environment to undergo subsequent mergers: the host’s escape velocity must exceed the gravitational recoil (kick) velocity of each merger remnant. Hierarchical BH mergers impart substantially larger kicks than first-generation mergers (see the left panel of Figure 2), so only environments with sufficiently deep potential wells–such as nuclear star clusters or AGN disks–could keep the GW231123 progenitors bound after each merger. Second, the binary orbits must be hardened efficiently so that multiple mergers can occur within a reasonable timescale. In dense stellar systems, dynamical interactions (e.g., binary-single encounters) help harden binaries, but an even more efficient mechanism is likely needed for rapid, repeated mergers. Gas dynamical friction in AGN disks can rapidly shrink binary orbits and shorten merger timescales (Yang et al., 2019), and simulations suggest that this gas-assisted hardening may dominate the upper limit on BBH merger masses (Vaccaro et al., 2024; Xue et al., 2025). Therefore, among known environments, AGN disks are the more plausible sites for producing BBH mergers as massive as GW231123.
After our initial manuscript appeared online, we noted that several alternative scenarios for GW231123 have also been discussed. Bartos & Haiman (2025) showed that accretion is also possible to produce mergers with near-equal, high-masses, and strongly aligned high spin mergers like GW231123, particularly in binaries embedded in AGN disks and Population III remnants. Meanwhile, Croon et al. (2025) found that stellar evolution alone is sufficient to explain GW231123 without accretion (see also (Baumgarte & Shapiro, 2025)). The direct formation of GW231123 via Population III was also proposed by Liu et al. (2025); Popa & de Mink (2025), and the possible primordial black hole origins were investigated by De Luca et al. (2025); Yuan et al. (2025). Additionally, some other explanations for the signal of GW231123 were also suggested (see e.g., He et al., 2025; Liu & Lai, 2025; Ray et al., 2025; Chakraborty & Mukherjee, 2025; Yang et al., 2025; Hu et al., 2025; Bini et al., 2026).
@@
If AGN-disk channels indeed dominate the formation of GW231123-like events, then BBH merger masses up to $\sim\mathcal{O}(10^{4})M_{\odot}$ might be achievable (Vaccaro et al., 2024), which are well beyond the detection range of the current generation of ground-based detectors (LIGO Scientific Collaboration et al., 2015; Acernese et al., 2015; Akutsu et al., 2018). The next generation of gravitational-wave observatories, including planned space-based detectors (Luo et al., 2016; Hu & Wu, 2017; Amaro-Seoane et al., 2017; Abbott et al., 2017; Punturo et al., 2010), will be required to observe such extremely massive mergers, offering a more complete view of the BBH mass spectrum and evolutionary pathways. Beyond standard astrophysical channels, a few exotic formation scenarios have been proposed. For example, “dark stars” (hypothetical early-universe stars powered by dark matter annihilation) could form IMBHs; however, the expected masses of such remnants are far larger than those of GW231123 (Lei et al., 2025a). Another proposal involves a cosmological coupling mechanism that gradually increases BH masses, which in principle might produce BHs as massive as GW231123 (Croker et al., 2021). This cosmological coupling hypothesis, however, has been ruled out by very recent JWST observations (Lei et al., 2025b). Ultimately, a comprehensive population analysis of BBH mergers will be necessary to fully elucidate the origins and formation channels of GW231123 and similar exceptional events (e.g., Li et al., 2021, 2022; Franciolini & Pani, 2022; Abbott et al., 2023b; Li et al., 2024b; Guo et al., 2024; Alvarez-Lopez et al., 2025).
@@
This work is supported by the National Natural Science Foundation of China (No. 12233011, No. 12503059, and No. 12303056), the General Fund (No. 2024M753495) of the China Postdoctoral Science Foundation, and the Priority Research Program of the Chinese Academy of Sciences (No. XDB0550400). This research has made use of data and software obtained from the Gravitational Wave Open Science Center (https://www.gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration and the Virgo Collaboration. LIGO is funded by the U.S. National Science Foundation. Virgo is funded by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale della Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by Polish and Hungarian institutes.
Appendix A Reweighted by Population-informed Priors
In hierarchical Bayesian inference, combining multiple events within a population allows the prior for individual event parameters to be informed by the learned population distribution (Fishbach et al., 2020; Miller et al., 2020). This approach enables the self-consistent joint inference of both the population shape, such as mass and spin distributions, and the parameters of individual events (Mandel, 2010). A comprehensive and rigorous approach, requires a population analysis that fits for the parameters of GW231123 together with full BBH observations (see Callister, 2021, for a technical guidance). However, at the time of the initial submission of this paper, the data from GWTC-4 had not been fully released  [^8] Please refer Li et al. (2025b); Wang et al. (2025b) for our consequent analysis when the full GWTC-4 data was released. The distribution of the populations are broadly consistent with previous for GWTC-3(Li et al., 2024b), except for the mass range of high-spin subpopulation. The difference is likely caused by the prior range of maximum mass, which was assumed to be $100M_{\odot}$ in GWTC-3.. Therefore, in this work we employ a simpler and less rigorous approach (see also Fishbach & Holz, 2020, for analysis of GW190521), and adopt the population-informed prior that is given by the posterior population distribution inferred with GWTC-3 (Li et al., 2024b),
$P_{\mathcal{H}}(\lambda|d_{\rm GWTC-3})\propto\int{p_{\mathcal{H}}(\lambda|\Lambda)p(\Lambda|d_{\rm GWTC-3}){\rm d}\Lambda},$ (A1)
To obtain the parameters of GW231123 ($\lambda_{0}$) under these priors, we reweight the posterior samples of LVK (Abac et al., 2025) by
$\frac{P_{\mathcal{H}}(\lambda_{0}|d_{\rm GWTC-3})}{p_{\rm default}(\lambda_{0})},$ (A2)
where $p_{\rm default}(\lambda_{0})$} is the default prior for parameter estimation of GW231123 (Abac et al., 2025). We consider two different scenarios to construct population-informed priors. One ($P_{\rm HS-HS}(\chi_{1},\chi_{2}|d_{\rm GWTC-3})$) is that both components of GW231123 belong to the high-spin (HS) subpopulation, the other ($P_{\rm HS-LS}(\chi_{1},\chi_{2}|d_{\rm GWTC-3})$) is that primary / secondary BH belong to HS / low-spin (LS) subpopulation. The scenarios that the primary BH belong to the LS subpopulation are not considered, because the $\chi_{1}$ posterior samples of GW231123 are scarcely supported in distribution of HS subpopulation, which is also reflected by the Bayes factors between different priors (see Table 1). In principle, GW231123 could have been directly reweighted using the predictive distribution of the two-component population (i.e., the mixture of low-spin (LS) and high-spin (HS) components) rather than by manually assigning each black hole to a specific component. We did not adopt this approach because the two-component model in Li et al. (2024b) was constructed in the $m-\chi$ plane and its applicable mass range, as inferred from GWTC-3, is limited to $100M_{\odot}$. Instead, for GW231123 we adopted a uniform prior distribution for $m_{1}$ and $m_{2}$, similar to the default prior used in parameter estimation (Abac et al., 2025). In both scenarios, the priors for spin tilt angles $P(\cos\theta_{1},\cos\theta_{2}|d_{\rm GWTC-3})$ are the same as that inferred in Li et al. (2024b).
Figure 4 shows the reweighed posterior samples compared to the samples under default prior. We find that under HS-HS prior, the $m_{2}$ shows more / less support for the high-mass / low-mass peak. Whereas $m_{2}$ shows more support for the peak at $\sim 60M_{\odot}$ under the HS-LS prior. If the secondary of GW231123 was a $\sim 60M_{\odot}$ and slowly spinning BH, it would support the existence of a group of low-spin $\sim 50-70M_{\odot}$ BHs (Wang et al., 2025b). However the HS-LS prior is less favored compared to the HS-HS prior, as show in Table 1.
Figure 4: Posterior distributions of component masses, spin magnitudes, and spin orientations of GW231123 under different priors.
Appendix B Bayes Factors between population-informed Priors
In this section, we perform a systematic comparison of the different spin priors discussed in the main text via the computation of their Bayes factors using GW231123 data. The aim of this section is not to identify the parameter distribution that best fits the data – such an endeavor would require a joint analysis of multiple BBH events – but rather to evaluate the plausibility of a given prior assumption $\mathcal{H}$: GW231123 belong to one of the inferred subpopulation from GWTC-3, $p_{\mathcal{H}}(\lambda|d_{\rm GWTC-3})$, in the context of the single-event likelihood $p(d_{\rm GW231123}|\lambda)$ (see also Fishbach & Holz, 2020, for GW190521). Then the Bayesian evidence for $\mathcal{H}$ given data $d_{\rm GW231123}$ is
$p(d_{\rm GW231123}|\mathcal{H},d_{\rm GWTC-3})=\int{p(d_{\rm GW231123}|\lambda)p_{\mathcal{H}}(\lambda|\Lambda){\rm d}\lambda p(\Lambda|d_{\rm GWTC-3}){\rm d}\Lambda},$ (B1)
We calculate Eq.(B1) by Monte Carlo integration (or importance sampling):
$p(d_{\rm GW231123}|\mathcal{H},d_{\rm GWTC-3})=\left\langle\left\langle\frac{p_{\mathcal{H}}(\lambda|\Lambda)}{p_{\rm default}(\lambda)}\right\rangle_{\{\lambda\}}\right\rangle_{\{\Lambda\}|d_{\rm GWTC-3}}.$ (B2)
The Bayes factor of $\mathcal{H}_{A}$ over $\mathcal{H}_{B}$ is $\mathcal{B}^{A}_{B}=\frac{p(d_{\rm GW231123}|\mathcal{H}_{A},d_{\rm GWTC-3})}{p(d_{\rm GW231123}|\mathcal{H}_{B},d_{\rm GWTC-3})}$. The Bayes factors between hypotheses that the two components of GW231123 belong to LS or HS subpopulation are summarized in Table 1. We find that both BHs of GW231123 belong to the HS subpopulation is most favored, and the hypothesis that the primary BH belongs to the LS subpopulation is ruled out.
Table 1: Bayes factors of hypothesis $\mathcal{H}_{x}$ relative to the hypothesis that both components belong to the high-spin subpopulation $\mathcal{H}_{\rm HS-HS}$.
$\log_{10}\mathcal{B}_{\rm HS-HS}^{x}$ BH1 Spin magnitude
BH2 Spin magnitude $\chi_{1}\sim P_{\rm LS}(\chi|d_{\rm GWTC-3})$ $\chi_{1}\sim P_{\rm HS}(\chi|d_{\rm GWTC-3})$
$\chi_{2}\sim P_{\rm LS}(\chi|d_{\rm GWTC-3})$ 2.7-2.7 0.8{-0.8}
$\chi_{2}\sim P_{\rm HS}(\chi|d_{\rm GWTC-3})$ 2.7-2.7 $0.0$
Appendix C Mass function of higher-generation BHs
C.1 Details about mass function of different generations of BHs